Circular No. 7467

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 S4 (LINEAR) I. de Pater, University of California, Berkeley, with M. R. Hogerheijde,

M. C. H. Wright, R. Forster, W. Hoffman, L. E. Snyder, A. Remijan, L. M. Woodney, M. F. A’Hearn, P. Palmer, Y.-J. Kuan, H.-C. Huang, G. A. Blake, C. Qi, J. Kessler, and S.-Y. Liu, report the detection of HCN from comet C/1999 S4 at the Berkeley-Illinois-Maryland-Association (BIMA) Array in autocorrelation mode: “The peak antenna temperature in a 130′′ beam, averaged over July 21–24 (∼ 10 hr total on source), was 3.5 ± 1 mK (a signal suggesting an outgassing rate a few percent that of C/1995 O1). No signal was detected in cross-correlation mode with the combined Owens Valley Radio Observatory and BIMA data (the virtual Combined Array for Research in Millimeter-wave Astronomy). This suggests a source size slightly larger than expected for a Haser model.”

M. Kidger, Instituto de Astrofísica de Canarias, writes that nightly observations made since July 23 in U, B, V, R, and Z broadband filters with the 1-m Jacobus Kapteyn Telescope show what appears to be the complete disruption of the comet’s nucleus: “The central condensation was highly condensed and showed the typical ‘teardrop’ form on July 23.9 and 24.9 UT, although its brightness decreased by a factor of ∼ 3 between the two nights. On July 25.9 the central condensation was seen to be strongly elongated (length ∼ 15′′) in p.a. 80^{o}, with a very flat brightness distribution. The condensation’s brightness faded further and its length increased to ∼ 30′′ and 45′′–50′′ (p.a. 80^{o}) on July 26.9 and 27.9, respectively. On July 27.9, there was no evidence of any local brightness peak that would indicate the presence of subnuclei. The expansion velocity of the condensation is ∼ 40 m/s, indicating that it is particulate material and not gas. The gas tail, which virtually disappeared between July 23.9 and 24.9, has reformed as an extension of the major axis of the central condensation.”

COMET P/2000 O2 (KUSHIDA) C. E. Delahodde, European Southern Observatory, reports the recovery

by O. R. Hainaut and herself of comet P/1994 A1 (= 1994a = 1993 XX) with the 3.6-m reflector. The indicated correction to the prediction by S. Nakano on MPC 31664 is ΔT = −0.10 day. Further observations and improved orbital elements are given on MPEC 2000-O32.

2000 UT  α_{2000}  δ_{2000}
July 25.33038  23^{h}34^{m}26.^{s}95  +2^{o}52′51′′.7
27.24488  23 34 08.12  +2 53 42.6
2000 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green